OpenStudy (rissyroo118):

I can buy 56 diapers at Babies-R-Us for $28, and I can buy 112 diapers online for $52. Which is the better deal, and how much will you save per diaper? (Points : 5) Babies R Us by $0.04 Online by $0.04 Babies R Us by $0.15 Online by $0.15

OpenStudy (anonymous):

The better deal is 112 diapers for $52 and you would save $4

OpenStudy (gorv):

calculate value of one diaper for both purchase

OpenStudy (gorv):

at babies r u i diaper= 28/56=1/2=0.5

OpenStudy (gorv):

and online one diaper=52/112=0.46

OpenStudy (gorv):

so online is better

OpenStudy (gorv):

saving per diaper =0.5-0.46=0.04 dollar